Suppose 3.6 kg of ice at -4˚C is added to an aquarium containing 20 gallons of water initially at 25˚C. What is temperature of the water when all the ice melts and the system is at a constant temperature?
Answer #1

Mw = mass of water = 20 gallons = 75.71 kg

m = mass of ice = 3.6 kg

L = latent heat of fusion of ice to water = 334000 J/kg

cice = specific heat of ice = 2050

cwater = specific heat of water = 4186

T = final equilibrium temperature

using conservation of energy

m cice (4) + mL + m c (T) = M cwater (25 - T)

(3.6) (2050) (4) + 3.6 x 334000 + (3.6) (4186) (T) = (75.71) (4186) (25 - T)

(3.6) (2050) (4) + 3.6 x 334000 + (3.6) (4186) (T) = (75.71) (4186) (25 - T)

T = 20.2